# Unlocking the Potential of Model Calibration in Federated Learning

### v2

@@ 5.2 Experimental Results
Calibration strategy validation. Our calibration strategy applies a higher auxiliary penalty loss to local models that significantly deviate from the received global model, addressing global calibration needs during local training. To validate the soundness of our design, we experiment with an opposite logic, where a higher penalty is applied to local clients more similar to the received global model. We specifically adjust the weight $\beta_{m}$ for each client $m$ to $\texttt{sim}(\delta^{(t-1)},\delta_{m}^{(t,e)})^{-1}$. We conducted this destructive experiment to assess whether this approach affects FL and to validate our original calibration logic. In Table 3, we applied a reverse penalty strategy to optimal configurations of our method (i.e., NUCFL (DCA+L-CKA) and (NDCA+L-CKA)) on the CIFAR-100 dataset. The results show that using reversed logic negatively impacts the system, leading to a calibration penalty biased towards local heterogeneity and ignoring global calibration needs, thus degrading the FL system’s performance. This even resulted in poorer performance compared to uncalibrated FL as shown in Table 1. This experiment confirms the soundness of our design logic, which strategically prioritizes global calibration needs during local training and suggests the benefits of adhering to our approach to avoid disadvantageous effects.
Figure 3: Reliability diagrams for non-IID FedAvg ($\alpha=0.5$) with different calibration methods using the CIFAR-10 dataset. The lower ECE and smaller gap show the effectiveness of our method.
Figure 4: Comparison of confidence calibration across different FL settings using the MNIST dataset. The default setting is 50 clients, $\alpha=0.5$, 100% participation, with one variable changed at a time.
Calibration Method Non-IID FedAvg ($\alpha=0.5$)
@@
Uncal. 72.77 9.80 3.21
FedCal Peng et al. [2024] 72.77 5.97 3.03
NUCFL (DCA+L-CKA) 73.73 5.14 2.87
NUCFL (MDCA+L-CKA) 73.64 5.20 2.96
Table 4: Comparison with SOTA baseline.
Comparison with the existing model calibration baseline for FL. We now conduct comparisons with the only existing model calibration framework for FL, FedCal Peng et al. [2024]. Table 4 displays the calibration results of our method alongside the baseline using FedAvg under non-IID conditions ($\alpha=0.5$) on CIFAR-100 datatset. For the scaler used in their work, we adhered to their design, employing an MLP with layer sizes of 100-64-64-100. We see that our proposed NUCFL has shown superior performance compared to FedCal. While the use of a scaler aggregated solely from local clients can help reduce global ECE, it often overlooks the essential interaction between global and local calibration needs. Our method addresses this by considering these interactions, thus avoiding biases toward local heterogeneity and better accommodating the broader global calibration requirements, a challenge that the aggregation methods might not fully resolve. More comparisons using other datasets are available in Appendix B.4.
Calibration Method Non-IID FedAvg ($\alpha=0.5$)
@@
Uncal. 72.77 9.80 3.21
DCA Liang et al. [2020] 72.39 6.95 3.07
MDCA Hebbalaguppe et al. [2022b] 71.47 7.69 3.19
NUCFL (DCA+L-CKA) 73.73 5.14 2.87
TS Guo et al. [2017] 72.77 7.28 3.26
DCA+TS 72.39 6.10 2.83
MDCA+TS 71.47 7.33 3.21
NUCFL (DCA+L-CKA)+TS 73.73 4.98 2.65
Table 5: Comparison with post-hoc calibration.
Compatibility with post-hoc calibration. Temperature scaling (TS) Guo et al. [2017] is one of the most effective post-hoc calibration methods, which adjusts a model’s logits by a scaling factor to calibrate confidence scores without affecting accuracy. Like all post-hoc methods, TS requires a hold-off dataset to tune the model and learn the calibration parameter (i.e., temperature). However, this need conflicts with the data privacy policies of FL, making TS impractical in such settings. For this analysis, we set aside these privacy concerns to compare with TS, assuming that data from each client is available at the server. Under these assumptions, we conduct non-IID ($\alpha=0.5$) FedAvg and sample five images from each client to create a server dataset for post-hoc calibration. Table 5 compares train-time and post-hoc calibration methods using the CIFAR-100 dataset. Results show TS outperforming some train-time methods, like MDCA, while our method still provides superiority in both accuracy and calibration error. Recall from Remark 2 that our NUCFL can also be complemented by post-hoc methods when a holdout dataset is available for refining the global model. We apply post-hoc calibration on the global model using TS after train-time calibration during local training. This combined approach shows improved performance over train-time calibration alone. Notably, integrating TS with NUCFL yields the best results, significantly reducing error.
Further experiments. Additional results on varying degrees of non-IIDness, evaluations across datasets, and discussions on how our method tackles the over/under-confidence problem are available in Appendix B.
@@ 6 Conclusion
We conducted one of the earliest systematic study on model calibration for FL and introduced a novel framework, NUCFL, designed to ensure a well-calibrated global model. NUCFL measures the similarity between client and server models and dynamically adjusts calibration penalties for each client, addressing both heterogeneity and global calibration needs effectively. Through extensive experiments, NUCFL demonstrated significant improvements over existing calibration methods, proving its adaptability with various types of auxiliary calibration penalties and compatibility across different FL algorithms. Our research seeks to deepen understanding of the underlying effects of model calibration on FL, enhancing the trustworthiness of FL systems.
@@ A.1 Calibration Baselines
We offer an overview of each compared calibration method, including the hyperparameter settings utilized during training. Focal loss Lin et al. [2017] is defined as $\text{FL}=-\sum^{N}_{i=1}(1-p_{i,y_{i}})^{\gamma}\cdot\log(p_{i,y_{i}})$, where $p_{i,y_{i}}$ is the predicted confidence score for sample $i$ and $\gamma$ is a hyper-parameter. We trained models with $\gamma\in\{1,2\}$. Label Smoothing (LS) Müller et al. [2019] is defined as $\text{LS}=-\sum^{N}_{i=1}\sum^{K}_{j=1}q_{i,j}\cdot\log(p_{i,j})$, where $p_{i,j}$ is the predicted confidence score for class $j$ of sample $i$, and $q_{i}$ is the soft target vector for sample $i$, such that $q_{i,j}=\frac{\alpha}{K-1}$ if $j\neq y_{i}$, else $q_{i,j}=(1-\alpha)$. $\alpha$ is a hyperparameter set at 0.1, following established literature settings Hebbalaguppe et al. [2022b]. For the Brier Score (BS) Brier [1950], we train using a loss defined as the squared difference between the predicted probability vector and the one-hot target vector. For Maximum Mean Calibration Error (MMCE) Kumar et al. [2018], we use the weighted MMCE, with $\lambda$ selected from $\{2,4\}$, as a regularization term along with cross-entropy. For FLSD Mukhoti et al. [2020], we use $\gamma=3$, reported as the best configuration in their paper. For Margin-based Label Smoothing (MbLS) Liu et al. [2021], we set the margin at 10 and $\mu=0.1$, configurations reported as optimal in their paper. For DCA Liang et al. [2020], we train using a loss composed of cross-entropy plus $\beta\cdot\ell_{dca}$, where $\beta$ is a hyperparameter selected from $\{1,5,10\}$ as demonstrated in their paper. Similarly, for MDCA Hebbalaguppe et al. [2022b], we apply the cross-entropy loss enhanced by $\beta\cdot\ell_{mdca}$, with $\beta$ also chosen from $\{1,5,10\}$ as well. For all baselines, we report results that show the lowest ECE among their hyperparameter settings in our paper.
@@ A.2 Federated Algorithms
For all FL algorithms, we distributed the training data across $M=50$ clients and performed 100 rounds of training. Each client conducted 5 local training epochs per round. The batch size was set at 64, and we explored learning rates within the range [1e-2, 5e-3, 1e-3, 5e-4]. For FedAvg McMahan et al. [2017], clients perform standard stochastic gradient descent (SGD) updates as ClientOPT and upload their gradients to the central server. On the server side, the global model is updated using SGD as ServerOPT with the received gradients from the clients. FedProx Sahu et al. [2018] introduces a proximal term during local client training as their ClientOPT, to address data heterogeneity issues by preventing significant drifts in client model updates. It employs a weighted average for aggregation and updates the global model using SGD, similar to ServerOPT in FedAvg. We set the proximal term coefficient parameter $\mu=1$ for the proximal term coefficient in FedProx. Scaffold Karimireddy et al. [2019] employs client-variance reduction as its ClientOPT and corrects model drifts in local updates by introducing control variates. For ServerOPT, Scaffold incorporates the clients’ gradients and control variates for aggregation, and updates the global model using SGD. FedDyn Acar et al. [2021] dynamically adjusts the local objectives by incorporating two regularizers as ClientOPT to ensure that the local optimum aligns asymptotically with the stationary points of the global objective. We set the parameter $\alpha=0.1$ as commonly used in the literature, and the local model is updated using SGD. Additionally, they introduce their server update as ServerOPT through the use of server state vectors. FedNova Wang et al. [2020b] uses normalized stochastic gradients as ClientOPT to compensate for data imbalances across clients, and then introduces a weighted rescaling process as their ServerOPT. We run all experiments on a 3-GPU cluster of Tesla V100 GPUs, with each GPU having 32GB of memory.
@@ B.1 Calibration on Centralized Learning
We present train-time calibration methods for traditional centralized training in Table 6. We train AlexNet using the SGD optimizer on the entire CIFAR-100 training dataset for 100 epochs with various train-time calibration methods. We found that most of these methods effectively calibrate in centralized settings, maintaining accuracy and exhibiting lower calibration errors.
Calibration Method Centralized Training
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 76.53 4.338 2.597
Focal Lin et al. [2017] 75.70 4.079 2.240
LS Müller et al. [2019] 76.63 4.145 2.466
BS Brier [1950] 75.52 5.464 3.190
MMCE Kumar et al. [2018] 75.31 4.329 2.690
FLSD Mukhoti et al. [2020] 75.49 5.162 3.145
MbLS Liu et al. [2021] 76.57 4.011 2.428
DCA Liang et al. [2020] 77.24 3.950 2.167
MDCA Hebbalaguppe et al. [2022b] 76.69 3.856 2.040
Table 6: Performance of centeralized training with various calibration methods.
@@ B.2 Additional Results of Calibration on FL
This section explores the performance of each calibration method on different datasets and data distributions for various FL algorithms. To ensure consistency in our experiments, most of the configuration remains unchanged. Specifically, we set the number of clients ($M$) to 50 and the total number of federated rounds ($T$) to 100, with each client performing $E=5$ epochs of local training. Additionally, the hyperparameter search space and the details for each FL algorithm are consistent, as detailed in Appendix A. Tables 7,  8,  9, and 10 display the performance and calibration metrics of various FL algorithms on the MNIST dataset across IID and non-IID conditions, with $\alpha$ values of 0.5, 0.1, and 0.05, respectively. Similarly, Tables 11,  12,  13, and  14 present the performance and calibration metrics of different FL algorithms on the CIFAR-10 dataset under IID and non-IID conditions, with corresponding $\alpha$ values of 0.5, 0.1, and 0.05, respectively. Across these experimental results, which consider various data distribution setups and different datasets, our generic framework NUCFL consistently demonstrates superiority over the baseline by achieving lower calibration error. The performance improvements of NUCFL with different configurations, such as various similarity measurements and different auxiliary calibration losses, highlight the effectiveness and flexibility of our method.

### v3

@@ 5.2 Experimental Results
Calibration strategy validation. Our calibration strategy assigns higher auxiliary penalty loss to local models that deviate more from the global model, addressing global calibration needs during local training. To validate our design, we test an opposite logic, applying higher penalties to clients more similar to the global model. We adjust the weight $\beta_{m}$ for each client $m$ to $\texttt{sim}(\delta^{(t-1)},\delta_{m}^{(t,e)})^{-1}$. We conducted this experiment to assess whether reversing our calibration logic affects FL performance and to validate our original design. In Table 3, we applied a reverse penalty strategy to optimal configurations of our method (i.e., NUCFL (DCA/MDCA + L-CKA)) on the CIFAR-100 dataset. The results show that reversed logic degrades performance by biasing calibration towards local heterogeneity while neglecting global needs, even performing worse than uncalibrated FL (Table 1). This confirms the soundness of our design, which strategically prioritizes global calibration needs during local training.
Comparison with FedCal (Peng et al., 2024). We now conduct comparisons with the only existing model calibration framework for FL, FedCal (Peng et al., 2024). FedCal bases its algorithm on FedAvg; thus, we display the calibration results of our method alongside the baseline using FedAvg under non-IID conditions ($\alpha=0.5$) on CIFAR-100 datatset in Table 4.
Calibration Method Non-IID FedAvg
@@
Uncal. 61.34 10.52 3.61
FedCal (Peng et al., 2024) 61.34 8.80 3.49
NUCFL (DCA+L-CKA) 62.05 6.14 3.07
NUCFL (MDCA+RBF-CKA) 61.54 6.20 3.09
Table 4: Comparison with FedCal.
For the scaler used in their work, we adhered to their design, employing an MLP with layer sizes of 100-64-64-100. We see that our proposed NUCFL shows superior performance compared to FedCal. While using a scaler aggregated from local clients can reduce global ECE, it may neglects the interactions between global and local calibration needs. Our method considers these interactions, minimizing local biases and better meeting global calibration requirements. More comparisons using other datasets and data distribution are available in Appendix B.4. The superiority of our method in both accuracy and calibration error shows NUCFL’s effectiveness in complex FL environments.
Calibration Method Non-IID FedAvg
@@
Uncal. 61.34 10.52 3.61
DCA 61.24 7.69 3.24
MDCA 61.03 7.71 3.29
NUCFL (DCA+L-CKA) 62.05 6.14 3.07
NUCFL (MDCA+L-CKA) 62.17 6.25 3.11
TS (Guo et al., 2017) 61.34 8.22 3.43
DCA+TS 61.24 7.00 3.03
MDCA+TS 61.03 7.29 3.11
NUCFL (DCA+L-CKA)+TS 62.05 5.91 2.97
NUCFL (MDCA+L-CKA)+TS 62.17 6.03 3.00
Table 5: Comparison and integration with post-hoc calibration. The results demonstrate our method’s compatibility with post-hoc calibration.
Compatibility with post-hoc calibration. Temperature scaling (TS) (Guo et al., 2017) is one of the most effective post-hoc calibration methods, which adjusts a model’s logits by a scaling factor to calibrate confidence scores without affecting accuracy. Like all post-hoc methods, TS requires a hold-off dataset to tune the model and learn the calibration parameter (i.e., temperature). However, this need conflicts with the data privacy policies of FL, making TS impractical in such settings. For this analysis, we set aside these privacy concerns to compare with TS, assuming that data from each client is available at the server. Under these assumptions, we conduct non-IID ($\alpha=0.5$) FedAvg and sample five images from each client to create a server dataset for post-hoc calibration. Table 5 compares train-time and post-hoc calibration methods using the CIFAR-100 dataset. Results show TS outperforming some train-time methods, like MDCA, while our method still provides superiority in both accuracy and calibration error. Recall from Remark 2 that our NUCFL can also be complemented by post-hoc methods when a holdout dataset is available for refining the global model. We apply post-hoc calibration on the global model using TS after train-time calibration during local training. This combined approach shows improved performance over train-time calibration alone. Notably, integrating TS with NUCFL yields the best results, significantly reducing error.
@@ 6 Conclusion
We conducted one of the earliest systematic study on model calibration for FL and introduced a novel framework, NUCFL, designed to ensure a well-calibrated global model. NUCFL measures the similarity between client and server models and dynamically adjusts calibration penalties for each client, addressing both heterogeneity and global calibration needs effectively. Through extensive experiments, NUCFL demonstrated significant improvements over existing calibration methods, proving its adaptability with various types of auxiliary calibration penalties and compatibility across different FL algorithms. Our research seeks to deepen understanding of the underlying effects of model calibration on FL, enhancing the trustworthiness of standard FL systems. Future work could explore calibration for personalized FL settings with personalized layers, unsupervised FL settings with dynamically adjusted calibration weighting for domain adaptation, and the use of similarity-based weighting for other auxiliary losses to enhance FL.
@@ A.1 Calibration Baselines
We offer an overview of each compared calibration method, including the hyperparameter settings utilized during training. Focal loss (Lin et al., 2017) is defined as $\text{FL}=-\sum^{N}_{i=1}(1-p_{i,y_{i}})^{\gamma}\cdot\log(p_{i,y_{i}})$, where $p_{i,y_{i}}$ is the predicted confidence score for sample $i$ and $\gamma$ is a hyper-parameter. We trained models with $\gamma\in\{1,2\}$. Label Smoothing (LS) (Müller et al., 2019) is defined as $\text{LS}=-\sum^{N}_{i=1}\sum^{K}_{j=1}q_{i,j}\cdot\log(p_{i,j})$, where $p_{i,j}$ is the predicted confidence score for class $j$ of sample $i$, and $q_{i}$ is the soft target vector for sample $i$, such that $q_{i,j}=\frac{\alpha}{K-1}$ if $j\neq y_{i}$, else $q_{i,j}=(1-\alpha)$. $\alpha$ is a hyperparameter set at 0.1, following established literature settings (Hebbalaguppe et al., 2022b). For the Brier Score (BS) (Brier, 1950), we train using a loss defined as the squared difference between the predicted probability vector and the one-hot target vector. For Maximum Mean Calibration Error (MMCE) (Kumar et al., 2018), we use the weighted MMCE, with $\lambda$ selected from $\{2,4\}$, as a regularization term along with cross-entropy. For FLSD (Mukhoti et al., 2020), we use $\gamma=3$, reported as the best configuration in their paper. For Margin-based Label Smoothing (MbLS) (Liu et al., 2021), we set the margin at 10 and $\mu=0.1$, configurations reported as optimal in their paper. For DCA (Liang et al., 2020), we train using a loss composed of cross-entropy plus $\beta\cdot\ell_{dca}$, where $\beta$ is a hyperparameter selected from $\{1,5,10\}$ as demonstrated in their paper. Similarly, for MDCA (Hebbalaguppe et al., 2022b), we apply the cross-entropy loss enhanced by $\beta\cdot\ell_{mdca}$, with $\beta$ also chosen from $\{1,5,10\}$ as well. For all baselines, we report results that show the lowest ECE among their hyperparameter settings in our paper.
@@ A.2 Federated Algorithms
For all FL algorithms, we distributed the training data across $M=50$ clients and performed 100 rounds of training. Each client conducted 5 local training epochs per round. The batch size was set at 64, and we explored learning rates within the range [1e-2, 5e-3, 1e-3, 5e-4]. For FedAvg (McMahan et al., 2017), clients perform standard stochastic gradient descent (SGD) updates as ClientOPT and upload their gradients to the central server. On the server side, the global model is updated using SGD as ServerOPT with the received gradients from the clients. FedProx (Sahu et al., 2018) introduces a proximal term during local client training as their ClientOPT, to address data heterogeneity issues by preventing significant drifts in client model updates. It employs a weighted average for aggregation and updates the global model using SGD, similar to ServerOPT in FedAvg. We set the proximal term coefficient parameter $\mu=1$ for the proximal term coefficient in FedProx. Scaffold (Karimireddy et al., 2019) employs client-variance reduction as its ClientOPT and corrects model drifts in local updates by introducing control variates. For ServerOPT, Scaffold incorporates the clients’ gradients and control variates for aggregation, and updates the global model using SGD. FedDyn (Acar et al., 2021) dynamically adjusts the local objectives by incorporating two regularizers as ClientOPT to ensure that the local optimum aligns asymptotically with the stationary points of the global objective. We set the parameter $\alpha=0.1$ as commonly used in the literature, and the local model is updated using SGD. Additionally, they introduce their server update as ServerOPT through the use of server state vectors. FedNova (Wang et al., 2020b) uses normalized stochastic gradients as ClientOPT to compensate for data imbalances across clients, and then introduces a weighted rescaling process as their ServerOPT. For applying NUCFL to FL algorthms, we provide a detail framework of our NUCFL in Algorithm 2. We run all experiments on a 3-GPU cluster of Tesla V100 GPUs, with each GPU having 32GB of memory.
@@ B.1 Calibration on Centralized Learning
We present train-time calibration methods for traditional centralized training in Table 6. We train ResNet-34 using the SGD optimizer on the entire CIFAR-100 training dataset for 100 epochs with various train-time calibration methods. We found that most of these methods effectively calibrate in centralized settings, maintaining accuracy and exhibiting lower calibration errors.
Calibration Method MNIST (CNN) FEMNIST (CNN) CIFAR-10 (AlexNet) CIFAR-100 (ResNet-34)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 97.13 0.74 0.26 94.29 3.07 1.58 84.19 5.53 2.55 68.33 5.80 3.00
Focal (Lin et al., 2017) 97.09 0.70 0.25 94.33 2.98 1.55 84.17 5.01 2.43 68.42 5.66 2.88
LS (Müller et al., 2019) 97.22 0.61 0.23 94.51 2.89 1.53 84.69 5.27 2.46 68.19 5.59 2.85
BS (Brier, 1950) 96.99 0.70 0.24 95.02 3.00 1.59 84.23 5.44 2.53 68.59 5.60 2.87
MMCE (Kumar et al., 2018) 97.23 0.55 0.20 94.31 2.71 1.50 84.33 4.87 2.35 68.69 5.44 2.55
FLSD (Mukhoti et al., 2020) 97.11 0.70 0.24 94.11 2.95 1.54 84.11 2.88 1.54 68.44 5.69 2.88
MbLS (Liu et al., 2021) 97.55 0.54 0.20 94.30 2.66 1.41 84.29 3.66 2.04 68.71 4.14 1.96
DCA (Liang et al., 2020) 97.29 0.52 0.20 94.33 2.59 1.39 84.20 3.71 2.05 68.45 4.29 1.98
MDCA (Hebbalaguppe et al., 2022b) 94.22 0.55 0.20 94.37 2.70 1.43 84.22 3.82 2.08 68.44 4.19 1.98
Table 6: Performance of centralized training with various calibration methods across different datasets. Train-time calibration methods demonstrate improvements in calibration error compared to the uncalibrated (Uncal.) model.
@@ B.2 Additional Results of Calibration on FL
This section explores the performance of each calibration method on different datasets and data distributions for various FL algorithms. To ensure consistency in our experiments, most of the configuration remains unchanged. Specifically, we set the number of clients ($M$) to 50 and the total number of federated rounds ($T$) to 100, with each client performing $E=5$ epochs of local training. Additionally, the hyperparameter search space and the details for each FL algorithm are consistent, as detailed in Appendix A. Tables 7,  8,  9, and 10 display the performance and calibration metrics of various FL algorithms on the MNIST dataset across IID and non-IID conditions, with $\alpha$ values of 0.5, 0.1, and 0.05, respectively. Tables 11,  12,  13, and 14 show the performance using FEMNIST dataset across IID and non-IID conditions. Similarly, Tables 15,  16,  17, and  18 present the performance and calibration metrics of different FL algorithms on the CIFAR-10 dataset under IID and non-IID conditions, with corresponding $\alpha$ values of 0.5, 0.1, and 0.05, respectively. Finally, Tables 19,  20,  21, and 22 show the performance using CIFAR-100 dataset across IID and non-IID conditions. Table 26 shows the calibration results with different local epoch numbers for non-IID ($\alpha=0.5$) FedAvg using CIFAR-100 dataset. Across these experimental results, which consider various data distribution setups and different datasets, our generic framework NUCFL consistently demonstrates superiority over the baseline by achieving lower calibration error. The performance improvements of NUCFL with different configurations, such as various similarity measurements and different auxiliary calibration losses, highlight the effectiveness and flexibility of our method.
In addition to the widely used FL algorithms introduced in Section 5, we conducted additional experiments using FedSpeed (Sun et al., 2023), FedSAM (Qu et al., 2022), FedMR (Hu et al., 2023), and FedCross (Hu et al., 2022) which are considered as the SOTA FL optimization methods. We conducted our experiments using the CIFAR-100 dataset under a non-IID ($\alpha=0.5$) data distribution, following a setup similar to that in Table 1. For the hyperparameter of FedSpeed, we selected $\lambda=0.1$ as the default value provided by the authors, and we selected $\alpha=0.99$ for FedCross. In Table 25, we see that in the absence of any calibration, these advanced FL algorithms do improve accuracy compared to FedAvg (ACC: 61.34; ECE: 10.52). However, they also introduce higher calibration errors. By applying NUCFL to these advanced FL algorithms, we find that our approach effectively reduces calibration error and even outperforms the SOTA calibration method in terms of calibration results. These results also demonstrate the adaptability of our method, showing that it can be applied to any FL algorithm to improve calibration while maintaining accuracy.